OpenStudy (anonymous):

Use two variables and two equations. Solve by any method you choose. Write your answer in the appropriate form. 195 people went on a bus tour. They used 7 vehicles. Some were cars and some were buses. If each car holds 5 people and each bus holds 45 people, how many of each vehicle did they need to take?

OpenStudy (nirmalnema):

195 persons 7 vehicles were used 5 by 1 car at a time 45 by 1 bus at a time let no bus be x therefore no of car be 7-x so 45(x)+5(7-x)=195 45x+35-5x=195 40x=160 x=4 means 4 buses and 3 cars were required 4*45+3*5=195